OpenStudy (anonymous):

1. A sphere has a volume of 288π ft3. Find the surface area of the sphere. 864π ft2 48π ft2 144π ft2 96π ft2

OpenStudy (anonymous):

please help

OpenStudy (ddcamp):

v = 4/3(pi)(r³) sa = 4(pi)(r²) v = sa * (1/3)(r) sa = 3*v*(1/r) r = cube root(3v/4pi)

OpenStudy (anonymous):

Thanks